The Supplement of an angle is 32 more than 2 times the complement of the angle. Find the measure of the angle, the complement, and the supplement.
Answer by nerdybill(5404)

The Supplement of an angle is 32 more than 2 times the complement of the angle. Find the measure of the angle, the complement, and the supplement.
.
The problem expect you to know that:
"angle" plus its supplement is 180.
"angle" plus its complement is 90.
.
Let x = the angle
then
180-x = supplement of the angle
90-x = complement of the angle
.
So, we have:
180-x = 2(90-x)+32
180 - x = 180 - 2x + 32
-x = -2x + 32
x = 32 degrees (the angle)
90-x = 90-32 = 58 degrees (complement)
180-x = 180-32 = 148 degrees (supplement)